# Generalised Fermat equation: a survey of solved cases

### v1

@@ 5.2 The general $a,b,c$
In this section we discuss cases of equation (1.2) with non-equal exponents $(p,q,r)$ such that at least two of the coefficients $a,b,c$ are greater than $1$.
In 2022, Chałupka et.al. [37] studied equations
@@
has no non-trivial primitive solutions.
@@
$p\equiv 3\text{ or }55\,\,(\text{mod }106)\text{\quad or \quad}p\equiv 47,65,113,139,143\,\,\text{or}\,\,167\,\,(\text{mod }168).$
Then the equation (5.7) has no non-trivial primitive solutions.
@@
If $a\in\{11,19,43,67,163\}$ and $p\geq 2$ is an integer, then the equations (5.5) and (5.6) have no non-trivial primitive integer solutions.
Noubissie [107] also considered equations (5.5) and (5.6) with $a=3$, and proved that for prime $p\geq 3$ these equations have no non-trivial primitive integer solutions with $6|y$.
Chałupka et.al. [37] also studied some equations (1.2) with $p=2q$. Specifically, they proved the following results.
@@
They also prove that for all primes $p\geq 11$ with $p\neq 19$, the equation (5.8) has no solutions in coprime odd integers $(x,y,z)$.
@@
Let $p$ be a prime satsifying
$p\equiv 3\text{ or }55\,\,(\text{mod }106)\text{\quad or \quad}p\equiv 47,65,113,139,143\text{ or }167\,\,(\text{mod }168).$
Then the equation (5.9) has no primitive solutions in odd integers.
@@
where $s$ is prime, and provided an algorithmically testable set of conditions on $(a,s)$ which, if satisfied, imply the existence of a constant $B=B(a,s)$ such that if $n>B$ then for any $k\geq 0$ equation (5.10) has no solutions in coprime integers $(x,y,z)$ such that $\gcd(ax,sy,z)=1$ and $z$ is even.
In a follow-on paper to [37], the authors [38] prove the partial result that for any prime $p=5$ or $11$ modulo $12$, the equation
@@
has no non-trivial primitive solutions $x,y,z$.
@@
$\cal{C}:=$$\{(11^{8},1,3^{6}\cdot 5^{10}),(3^{6}\cdot 11^{8},1,5^{10}),(3^{9}\cdot 11^{8},1,5^{10}),(1,11^{4},3^{6}\cdot 5^{10}),$ (5.12)
$(3^{4}\cdot 11^{8},1,5^{10}),(11^{8},3^{3},5^{10}),(3^{7}\cdot 11^{8},1,5^{10}),(3^{8}\cdot 11^{8},1,5^{10}),(3^{10}\cdot 11^{8},1,5^{10})\}.$
If $(a,b,c)\in\cal{C}$ then equation (5.11) has no non-trivial primitive integer solutions.
6 Summarising table
The below table summarises the equations solved in the literature considered in this survey. The table specifies the equation, any restrictions on the parameters, the progress towards the solution of the equation, and the reference to the solution. The progress we refer to can be:
@@ 6 Summarising table
Table 6.1: Generalised Fermat equations (1.2) with $11/p+1/q+1/r<$ 1 and $|abc|>1$ solved in the literature [^19] For equations (1.2) with $|abc|=1$, see Table 1.1.. Assume $n$ is an integer and $p,r,s$ are primes. Unless otherwise stated, assume $\alpha,\beta,\gamma\geq 0$.
@@
$x^{4}+3y^{4}=4z^{4}$ Solved [57, pp. 630]
$x^{4}+8y^{4}=9z^{4}$ Solved [57, pp. 630]
$ax^{4}+by^{4}=cz^{4}$ $|a|+|b|+|c|\leq 62$ Non-zero [75, Sec 6.2.2]
[143, Secs 6.4-5]
$ax^{4}+by^{4}=bz^{4}$ $|a|+2|b|\leq 38$ Non-trivial [143, Sec 6.15]
$x^{4}+by^{4}=z^{4}$ $1\leq b\leq 218$ Non-trivial [143, Sec 6.15]
@@
$x^{4}+2py^{4}=z^{4}$ Any of the conditions in Theorem 2.4 Solved [137]
$a^{3}x^{4}+b^{3}y^{4}=c^{3}z^{4}$ Conditions in Theorem 2.5 Solved [102, 103]
@@
$ax^{2}+y^{4}=z^{p}$ $(a=3,p>17)$ or $(a=5,p>499)$ or Solved [61, 111]
$(a=6,p>563)$ or $(a=7,p>349)$
$sx^{p}+y^{4}=z^{2}$ $s>3,s\equiv 3$ (mod $8$), $s\neq 2t^{2}+1$, Solved [55]
@@
$2^{\alpha}s^{\beta}x^{p}+y^{4}=z^{4}$ $\alpha\geq 0$, $\beta\geq 0$, $s\geq 3$, $s\neq 2^{2^{k}}+1$, Solved [55, 4]
@@
$2^{\alpha}s^{\beta}x^{p}+y^{4}=z^{4}$ $502\leq s<$ 50, $s\neq 5$, $s\neq 17$, $p>5$, Solved [4]
$\alpha,\beta\geq 0$
@@
$ax^{2m}+y^{2}=z^{n}$ Conditions in Theorem 2.17 Solved [93]
@@
$x^{2}+by^{6}=z^{p}$ $p\geq r_{0}$, $(b,r_{0})$ given by Table 2.1 Solved [95]
$x^{2}+b^{2}y^{2m}=z^{4n}$ $b>0$ odd with Partial [145]
prime factor $l=\pm 3$ (mod 8)
@@
$ax^{6}+by^{6}=bz^{6}$ $|a|+2|b|\leq 10$ Non-trivial [143, Sec 6.16]
$x^{14}+2^{\alpha}7^{\beta}y^{14}=z^{14}$ $\alpha\geq 0,\beta\geq 1$ Solved [57, pp. 736]
@@
$x^{p}+y^{p}=pz^{p}$ $p$ regular (see Theorem 3.4) Solved [57, pp. 759]
@@
$x^{p}+y^{p}=31^{\alpha}z^{p}$ $7\leq p\leq 10^{6}$, $\alpha\geq 0$ Solved [45, Sec 15.7]
@@
$x^{p}+y^{p}=s^{\alpha}z^{p}$ $3\leq s\leq 100$, $p\geq 5$, $\alpha\geq 0$, $s\neq 31$ Solved [120, 88],
[45, Thm 15.5.3]
$x^{p}+y^{p}=psz^{p}$ $p\geq 5$ regular, $s$ only divisible by primes Solved [130]
@@
$x^{p}+2^{\beta}y^{p}=s^{\alpha}z^{p}$ $s>3$, $s\neq 2^{n}\pm 1$, $\alpha,\beta\geq 0$, Solved [88]
$p>(1+\sqrt{8(s+1)})^{2(s-1)}$, $p\nmid\alpha$
$x^{p}+16y^{p}=s^{\alpha}z^{p}$ $\alpha\geq 0$, $s\geq 3$, $s\neq 17$, $p\neq s$, $p\geq 5$, Solved [88]
@@
$ax^{p}+by^{p}=16cz^{p}$ Prime factors of $abc$ are $1$ (mod 3), Solved [60]
@@
$ax^{p}+by^{p}=16cz^{p}$ $n>0$ not a divisor of 14,16 or 18, Solved [60]
Prime factors of $abc$ are ±$\pm 1$ (mod $n$),
@@
$ax^{p}+by^{p}=cz^{p}$ Conditions in Theorem 3.23 Solved [60]
$ax^{p}+by^{p}+2^{n}cz^{p}=0$ Any of the conditions in Theorem 3.24 Solved [60]
$ax^{p}+2^{r}by^{p}+2^{r}cz^{p}=0$ Any of the conditions in Theorem 3.24 Solved [60]
$ax^{n}+by^{n}=cz^{n}$ Conditions in Theorem 3.26 Solved [116, 118]
@@
$x^{3}+y^{3}=s^{\alpha}z^{r}$ $s=2^{a}3^{b}-1$, $a\geq 5$, $b\geq 1$, Solved [5]
$\alpha\geq 1$, $r>s^{2s}$,
$(\alpha,r,a,b)$ do not satisfy (4.6)
@@
$x^{3}+y^{3}=sz^{r}$ $(s,r)$ satisfy Table 4.1, $r\geq s^{2s}$ Solved [5]
@@
$x^{12}-y^{12}=cz^{n}$ $n\geq 3$, $c\geq 1$, $l$ prime divisors of $c$, Solved [91]
$\nu_{l}(c)<12$ and $l\not\equiv 1$ (mod 4)
@@
$\gamma$ divisible only by primes $\gamma_{i}\not\equiv 1$ (mod 13),
@@
$x^{14}+y^{14}=2^{\alpha}3^{\beta}5^{\gamma}cz^{r}$ $r\geq 17$, $\alpha\geq 2$ or $\beta>0$ or $\gamma>0$ Solved [68]
@@
$x^{5}+y^{5}=2^{\alpha}5^{\beta}s^{\gamma}z^{2}$ $s=\pm 3$ (mod 10), $\alpha\geq 0$, Solved [135, 136]
$\beta\geq 0$, $\gamma\geq 0$
$x^{n}+y^{n}=cz^{2}$ $c\in\{2,3,5,6,10,11,13,17\}$, $n\geq 5$ Solved [13]
@@
$x^{p}+y^{p}=2^{\alpha}cz^{2}$ $c$ odd square-free, $3\nmid c$, $7\nmid c$, Solved [55]
$p>c^{132c^{2}}$, $\alpha\geq 1$
@@
$c$ only divisible by primes $c_{i}\not\equiv 1$ (mod $p$),
@@
$x^{p}+y^{p}=cz^{3}$ $c\neq 0$ cube-free, $3|c$, $C=\mathrm{rad}(c)$, Solved [92]
$p>C^{10C^{2}}$
@@
$x^{p}+2^{\alpha}y^{p}=z^{2}$ $\alpha\geq 2$, $p\geq 5$ Solved [81, 45, Sec 15.3]
@@
$x^{p}+s^{\alpha}y^{p}=z^{3}$ $s\in\{5,11,13,23,29,31,41,43,47,53,$ Solved [14]
59,61,67,71,79,83,97}59,61,67,71,79,83,97\}59 , 61 , 67 , 71 , 79 , 83 , 97 }, $\alpha\geq 1$, $p\geq 11$,
$p\nmid s^{2}-1$, $(s,p)\not\in\{(13,19),(29,11),$
$(43,13),(47,13),(59,11),(61,61),$
(67,73),(79,97),(97,13),(97,79)}(67,73),(79,97),(97,13),(97,79)\}( 67 , 73 ) , ( 79 , 97 ) , ( 97 , 13 ) , ( 97 , 79 ) }
$x^{4}+by^{4}=z^{5}$ $b\in\{3,11,19,43,67,163\}$ Solved [132, Thm 2.8]
@@
$x^{p}+2^{\alpha}y^{p}=cz^{2}$ $c\in\{3,5,7,11,13,15,17\}$, $\alpha\geq\alpha_{0}$, Solved [13]
$p>c$, $(c,\alpha,p)\neq(11,3,13)$,
where $(c,\alpha_{0})$ is given by (4.27)
$x^{p}+s^{\alpha}y^{p}=2z^{2}$ $p\geq 11$, $s\in\{5,11,13\}$, $s\neq p$, $\alpha\geq 0$ Solved [13]
@@
$x^{p}+2^{\alpha}y^{p}=sz^{2}$ $\alpha\geq 2$, $s>5$, $s=\pm 3$ (mod 8), $p>s^{8s^{2}}$ Solved [146]
$x^{p}+s^{\alpha}y^{p}=3^{\beta}z^{3}$ $s\not\in\{5,3t^{3}\pm 1,9t^{3}\pm 1:t\in\mathbb{N}\}$, Solved [14]
@@
$(s,p)\neq(7,13),(13,7)$, $3\nmid\beta$, $\alpha,\beta\geq 1$
$x^{p}+3^{\alpha}y^{p}=c^{\beta}z^{3}$ $c\in\{2,3,5,7,11,13,15,17,19\}$, Solved [14]
@@
$x^{p}+s^{\alpha}y^{p}=Mz^{3}$ Conditions of Theorem 4.48 Solved [92]
@@
$ax^{3}+by^{3}=z^{n}$ Conditions of Theorem 4.49 Partial [9, Thm 13.2]
@@
$ax^{p}+by^{p}=z^{2}$ $p\geq 11$, $ab\in\{2^{\alpha}11^{\beta},2^{\alpha}13^{\beta},2^{\alpha}19^{\beta},$ Solved [13]
$2^{\alpha}29^{\beta},2^{\alpha}43^{\beta},2^{\alpha}53^{\beta},2^{\alpha}59^{\beta},$
$2^{\alpha}61^{\beta},2^{\alpha}67^{\beta}\}$, $a,b$ coprime,
@@
$p\nmid ab$, $ab,p,\alpha,\beta$ not in Table 4.2
$ax^{p}+by^{p}=z^{2}$ $p\geq 11$, $ab\in\{2\cdot 19^{\beta},4\cdot 11^{\beta},4\cdot 19^{\beta},$ Solved [13]
$4\cdot 43^{\beta},4\cdot 59^{\beta},4\cdot 61^{\beta},4\cdot 67^{\beta}\}$ ,
$a,b$ coprime, $\beta\geq 1$,
$p\nmid ab$, $ab,p,\alpha,\beta$ not in Table 4.2
$ax^{p}+by^{p}=z^{2}$ $a,b$ coprime, $ab=2^{\alpha}s^{\beta}t^{\gamma}$, $\beta,\gamma\geq 0$, Solved [13]
$\alpha\geq 6$, $(s,t)\in\{(3,31)(\beta\geq 1),$
$(5,11)(\alpha\geq 7),(5,19),(5,23)(\beta\geq 1),$
$(7,19)(\gamma\geq 1),(11,13),(11,23)(\beta\geq 1),$
@@
$p\geq 11$, $p\nmid ab$, $ab,p,\alpha,\beta$ not in Table 4.3
$ax^{p}+by^{p}=z^{2}$ $ab=2^{\alpha}3^{\beta}$, $\gcd(a,b)=1$, $\alpha\geq 6$, Solved [149]
$\beta\geq 0$, $p\geq 7$, $p\nmid ab$
@@
$ax^{p}+by^{p}=cz^{2}$ $(a,b,c)\in\{(2^{n},l^{m},1),(2^{n}l^{m},1,1),$ Solved [84]
$(1,l^{m},2)\}$, conditions below (4.37)
@@
$x^{2}+by^{2p}=z^{3}$ $p\geq 7$, $\alpha\geq 6$, $\beta\geq 0$, $b\in\{2^{\alpha},4\cdot 3^{2\beta}\}$ Solved [149]
$x^{2}+2^{\alpha}3^{\beta}31^{\gamma}y^{2p}=z^{3}$ Conditions of Theorem 5.2 Solved [149]
$x^{2p}+by^{2}=z^{5}$ $p\geq 7$, $b$ positive odd, $5\nmid h(-b)$ Partial [147]
@@
$x^{2}+2^{\alpha}5^{\beta}11^{\gamma}y^{2p}=z^{5}$ $\alpha\geq 3$, $p>11$, $\beta\geq 1$, $\gamma\geq 0$ Solved [147]
$x^{2}+2^{\alpha}5^{\beta}19^{\gamma}y^{2p}=z^{5}$ $\alpha\geq 2$, $p\geq 11$, $p\neq 19$, $\beta\geq 1$, $\gamma\geq 0$ Solved [147]
@@
$p\equiv 47,65,113,139,143$ or $167$ (mod 168)
$ax^{2}+y^{2n}=4z^{3}$ $a\in\{11,19,43,67,163\}$, $n\geq 2$ Solved [37]
$x^{2}+ay^{2n}=4z^{3}$ $a\in\{11,19,43,67,163\}$, $n\geq 2$ Solved [37]

### v2

@@ 5.2 The general $a,b,c$
In this section we discuss cases of Eq. (1.1) with non-equal exponents $(p,q,r)$ such that at least two of the coefficients $a,b,c$ are greater than $1$.
In 2022, Chałupka et al. [38] studied equations
@@
has no non-trivial primitive integer solutions.
@@
$p\equiv 3\text{ or }55(\text{mod }106)\quad\text{ or }\quad p\equiv 47,65,113,139,143\text{ or }167(\text{mod }168).$
Then Eq. (5.7) has no non-trivial primitive integer solutions.
@@
If $a\in\{11,19,43,67,163\}$ and $p\geq 2$ is an integer, then Eqs. (5.5) and (5.6) have no non-trivial primitive integer solutions.
Noubissie [110] also considered Eqs. (5.5) and (5.6) with $a=3$, and proved that for prime $p\geq 3$ these equations have no non-trivial primitive integer solutions with $6|y$.
Chałupka et al. [38] also studied some Eqs. (1.1) with $p=2q$. Specifically, they proved the following results.
@@
They also prove that for all primes $p\geq 11$ with $p\neq 19$, Eq. (5.8) has no solutions in coprime odd integers $(x,y,z)$.
@@
Let $p$ be a prime satisfying
$p\equiv 3\text{ or }55(\text{mod }106)\quad\text{ or }\quad p\equiv 47,65,113,139,143\text{ or }167(\text{mod }168).$
Then Eq. (5.9) has no primitive solutions in odd integers.
@@
where $s$ is prime, and provided an algorithmically testable set of conditions on $(a,s)$ which, if satisfied, imply the existence of a constant $B=B(a,s)$ such that if $n>B$ then for any $k\geq 0$ Eq. (5.10) has no solutions in coprime integers $(x,y,z)$ such that $\gcd(ax,sy,z)=1$ and $z$ is even.
In a follow-on paper to [38], the authors [39] prove that for any prime $p=5$ or $11$ modulo $12$, the equation
@@
has no non-trivial primitive integer solutions $x,y,z$.
@@
[…]
$(1,11^{4},3^{6}\cdot 5^{10}),(3^{4}\cdot 11^{8},1,5^{10}),(11^{8},3^{3},5^{10}),$
$(3^{7}\cdot 11^{8},1,5^{10}),(3^{8}\cdot 11^{8},1,5^{10}),(3^{10}\cdot 11^{8},1,5^{10})\}.$
If $(a,b,c)\in\mathcal{C}$ then Eq. (5.11) has no non-trivial primitive integer solutions.
6 Summarizing table
Table 6.1 below summarizes the equations solved in the literature considered in this survey. The table lists Eqs. (1.1) with $|abc|>1$, see Table 1.1 for the case $|abc|=1$. Table 6.1 specifies the equation, any restrictions on the parameters, the progress towards the solution of the equation, and the reference to the solution. The progress we refer to can be:
@@ 6 Summarizing table
Table 6.1: Generalised Fermat Eqs. (1.1) with $11/p+1/q+1/r<$ 1 and $|abc|>1$ solved in the literature. Assume $m,n$ are integers and $p,r,s$ are primes. Unless otherwise stated, assume $\alpha,\beta,\gamma\geq 0$.
@@
$x^{4}+3y^{4}=4z^{4}$ Solved [59, pp. 630]
$x^{4}+8y^{4}=9z^{4}$ Solved [59, pp. 630]
$ax^{4}+by^{4}=cz^{4}$ $|a|+|b|+|c|\leq 62$ Non-zero [77, Sec 6.2.2]
[148, Sec 6.4-5]
$ax^{4}+by^{4}=bz^{4}$ $|a|+2|b|\leq 38$ Non-trivial [148, Sec 6.15]
$x^{4}+by^{4}=z^{4}$ $1\leq b\leq 218$ Non-trivial [148, Sec 6.15]
@@
$x^{4}+2py^{4}=z^{4}$ Any of the conditions Solved [142]
in Theorem 2.4
$a^{3}x^{4}+b^{3}y^{4}=c^{3}z^{4}$ Conditions in Solved [105, 106]
Theorem 2.5
@@
$ax^{2}+y^{4}=z^{p}$ $(a=3,p>17)$ or Solved [65, 114]
$(a=5,p>499)$ or
$(a=6,p>563)$ or
$(a=7,p>349)$
$sx^{p}+y^{4}=z^{2}$ $s>3,s\equiv 3$ (mod 8), Solved [57]
$s\neq 2t^{2}+1$,
@@
$2^{\alpha}s^{\beta}x^{p}+y^{4}=z^{4}$ $\alpha\geq 0$, $\beta\geq 0$, $s\geq 3$, Solved [57, 5]
$s\neq 2^{2^{k}}+1$,
@@
$2^{\alpha}s^{\beta}x^{p}+y^{4}=z^{4}$ $502\leq s<$ 50, $s\neq 5$, $s\neq 17$, Solved [5]
$p>5$, $\alpha,\beta\geq 0$
@@
$ax^{2m}+y^{2}=z^{n}$ Conditions in Solved [95]
Theorem 2.17
@@
$x^{2}+by^{6}=z^{p}$ $p\geq r_{0}$, $(b,r_{0})$ given by Solved [97]
Table 2.1
$x^{2}+b^{2}y^{2m}=z^{4n}$ $b>0$ odd with prime Partial [150]
factor $l=\pm 3$ (mod 8)
@@
$ax^{6}+by^{6}=bz^{6}$ $|a|+2|b|\leq 10$ Non-trivial [148, Sec 6.16]
$x^{14}+2^{\alpha}7^{\beta}y^{14}=z^{14}$ $\alpha\geq 0,\beta\geq 1$ Solved [59, pp. 736]
@@
$x^{p}+y^{p}=pz^{p}$ $p$ regular Solved [59, pp. 759]
(see Theorem 3.4)
@@
$x^{p}+y^{p}=31^{\alpha}z^{p}$ $7\leq p\leq 10^{6}$, $\alpha\geq 0$ Solved [47, Sec 15.7]
@@
$x^{p}+y^{p}=s^{\alpha}z^{p}$ $3\leq s\leq 100$, $p\geq 5$, Solved [124, 90],
$\alpha\geq 0$, $s\neq 31$ [47, Th.15.5.3]
$x^{p}+y^{p}=psz^{p}$ $p\geq 5$ regular, Solved [135]
$s$ only divisible by primes
@@
$x^{p}+2^{\beta}y^{p}=s^{\alpha}z^{p}$ $s>3$, $s\neq 2^{n}\pm 1$, Solved [90]
$p>(1+\sqrt{8(s+1)})^{2(s-1)}$,
$\alpha,\beta\geq 0$, $p\nmid\alpha$
$x^{p}+16y^{p}=s^{\alpha}z^{p}$ $\alpha\geq 0$, $s\geq 3$, $s\neq 17$, Solved [90]
$p\neq s$, $p\geq 5$,
@@
$ax^{p}+by^{p}=16cz^{p}$ Prime factors of $abc$ Solved [64]
are 1 (mod 3),
@@
$ax^{p}+by^{p}=16cz^{p}$ $n>0$ not a divisor of Solved [64]
14,16 or 18, prime factors
of $abc$ are ±$\pm 1$ (mod $n$),
@@
$ax^{p}+by^{p}=cz^{p}$ Conditions in Solved [64]
Theorem 3.23
$ax^{p}+by^{p}+2^{n}cz^{p}=0$ Any of the conditions Solved [64]
in Theorem 3.24
$ax^{p}+2^{r}by^{p}+2^{r}cz^{p}=0$ Any of the conditions Solved [64]
in Theorem 3.24
$ax^{n}+by^{n}=cz^{n}$ Conditions in Solved [120, 122]
Theorem 3.26
@@
$x^{3}+y^{3}=s^{\alpha}z^{r}$ $s=2^{a}3^{b}-1$, $a\geq 5$, Solved [6]
$b\geq 1$, $\alpha\geq 1$, $r>s^{2s}$,
$(\alpha,r,a,b)$ do not
satisfy (4.6)
@@
$x^{3}+y^{3}=sz^{r}$ $(s,r)$ satisfy Table 4.1, Solved [6]
$r\geq s^{2s}$
@@
$x^{12}-y^{12}=cz^{n}$ $n\geq 3$, $c\geq 1$, Solved [93]
$l$ prime divisors of $c$,
$\nu_{l}(c)<12$
and $l\not\equiv 1$ (mod 4)
@@
$\gamma$ divisible only by primes
$\gamma_{i}\not\equiv 1$ (mod 13),
@@
$x^{14}+y^{14}=2^{\alpha}3^{\beta}5^{\gamma}cz^{r}$ $r\geq 17$, $\alpha\geq 2$ or Solved [70]
$\beta>0$ or $\gamma>0$
@@
$x^{5}+y^{5}=2^{\alpha}5^{\beta}s^{\gamma}z^{2}$ $s=\pm 3$ (mod 10), Solved [140, 141]
$\alpha\geq 0$, $\beta\geq 0$, $\gamma\geq 0$
$x^{n}+y^{n}=cz^{2}$ $c\in\{2,3,5,6,10,$ Solved [15]
11,13,17}11,13,17\}11 , 13 , 17 }, $n\geq 5$
@@
$x^{p}+y^{p}=2^{\alpha}cz^{2}$ $c$ odd square-free, $3\nmid c$, Solved [57]
$7\nmid c$, $p>c^{132c^{2}}$, $\alpha\geq 1$
@@
$c$ only divisible by primes
$c_{i}\not\equiv 1$ (mod $p$),
@@
$x^{p}+y^{p}=cz^{3}$ $c\neq 0$ cube-free, $3|c$, Solved [94]
$C=\mathrm{rad}(c)$, $p>C^{10C^{2}}$
@@
$x^{p}+2^{\alpha}y^{p}=z^{2}$ $\alpha\geq 2$, $p\geq 5$ Solved [83]
[47, Sec 15.3]
@@
$x^{p}+s^{\alpha}y^{p}=z^{3}$ $s\in\{5,11,13,23,29,31,$ Solved [16]
$41,43,47,53,59,61,67,$
71,79,83,97}71,79,83,97\}71 , 79 , 83 , 97 }
$\alpha\geq 1$, $p\geq 11$, $p\nmid s^{2}-1$,
$(s,p)\not\in\{(13,19),(29,11),$
$(43,13),(47,13),(59,11),$
$(61,61),(67,73),(79,97),$
(97,13),(97,79)}(97,13),(97,79)\}( 97 , 13 ) , ( 97 , 79 ) }
$x^{4}+by^{4}=z^{5}$ $b\in\{3,11,19,43,67,163\}$ Solved [137, Th.2.8]
The case $a=1$ and fixed $p$
$x^{5}+7y^{5}=2^{\alpha}5^{\beta}z^{r}$ $r>41,\alpha\in\{1,2,3,4\},\beta\in\{2,3,4\}$ Solved [2]
$x^{5}+7y^{5}=z^{r}$ $r>41$ Partial [2]
@@
$x^{p}+2^{\alpha}y^{p}=cz^{2}$ $c\in\{3,5,7,11,13,15,17\}$, Solved [15]
$\alpha\geq\alpha_{0}$, $p>c$,
$(c,\alpha,p)\neq(11,3,13)$,
where $(c,\alpha_{0})$
is given by (4.25)
$x^{p}+s^{\alpha}y^{p}=2z^{2}$ $p\geq 11$, $s\in\{5,11,13\}$, Solved [15]
$s\neq p$, $\alpha\geq 0$
@@
$x^{p}+2^{\alpha}y^{p}=sz^{2}$ $\alpha\geq 2$, $s>5$, Solved [151]
$s=\pm 3$ (mod 8),
$p>s^{8s^{2}}$
$x^{p}+s^{\alpha}y^{p}=3^{\beta}z^{3}$ $s\not\in\{5,3t^{3}\pm 1,$ Solved [16]
$9t^{3}\pm 1:t\in\mathbb{N}\}$
@@
$(s,p)\neq(7,13),(13,7)$,
$3\nmid\beta$, $\alpha,\beta\geq 1$
$x^{p}+3^{\alpha}y^{p}=c^{\beta}z^{3}$ $c\in\{2,3,5,7,11,13,$ Solved [16]
15,17,19}15,17,19\}15 , 17 , 19 },
@@
$x^{p}+s^{\alpha}y^{p}=Mz^{3}$ Conditions of Solved [94]
Theorem 4.48
@@
$ax^{3}+by^{3}=z^{n}$ Conditions of Partial [10, Th.13.2]
Theorem 4.49
@@
$ax^{p}+by^{p}=z^{2}$ $p\geq 11$, $ab\in\{2^{\alpha}11^{\beta},2^{\alpha}13^{\beta},$ Solved [15]
$2^{\alpha}19^{\beta},2^{\alpha}29^{\beta},2^{\alpha}43^{\beta},$
$2^{\alpha}53^{\beta},2^{\alpha}59^{\beta},2^{\alpha}61^{\beta},2^{\alpha}67^{\beta}\}$,
$a,b$ coprime, $p\nmid ab$,
@@
$ab,p,\alpha,\beta$ not in Table 4.2
$ax^{p}+by^{p}=z^{2}$ $p\geq 11$, $ab\in\{2\cdot 19^{\beta},4\cdot 11^{\beta},$ Solved [15]
$4\cdot 19^{\beta},4\cdot 43^{\beta},4\cdot 59^{\beta},$
$4\cdot 61^{\beta},4\cdot 67^{\beta}\},$
$a,b$ coprime, $p\nmid ab$, $\beta\geq 1$,
$ab,p,\alpha,\beta$ not in Table 4.2
$ax^{p}+by^{p}=z^{2}$ $a,b$ coprime, $ab=2^{\alpha}s^{\beta}t^{\gamma}$, Solved [15]
$\beta,\gamma\geq 0$, $\alpha\geq 6$,
$(s,t)\in\{(3,31)(\beta\geq 1),$
$(5,11)(\alpha\geq 7),(5,19),$
$(5,23)(\beta\geq 1),(7,19)(\gamma\geq 1),$
$(11,13),(11,23)(\beta\geq 1),$
@@
$p\geq 11$, $p\nmid ab$,
$ab,p,\alpha,\beta$ not in Table 4.3
$ax^{p}+by^{p}=z^{2}$ $ab=2^{\alpha}3^{\beta}$, $\gcd(a,b)=1$, Solved [154]
$\alpha\geq 6$, $\beta\geq 0$, $p\geq 7$, $p\nmid ab$
@@
$ax^{p}+by^{p}=cz^{2}$ $(a,b,c)\in\{(2^{n},l^{m},1),$ Solved [86]
$(2^{n}l^{m},1,1),(1,l^{m},2)\}$,
conditions below (4.35)
@@
$7x^{p}+y^{p}=3z^{5}$ $p>71$ Partial [2]
$7x^{p}+2^{\alpha}5^{\beta}y^{p}=3z^{5}$ $p>71,\alpha\in\{1,2,3,4\},\beta\in\{3,4\}$ Solved [2]
@@
$x^{2}+by^{2p}=z^{3}$ $p\geq 7$, $\alpha\geq 6$, $\beta\geq 0$, Solved [154]
$b\in\{2^{\alpha},4\cdot 3^{2\beta}\}$
$x^{2}+2^{\alpha}3^{\beta}31^{\gamma}y^{2p}=z^{3}$ Conditions of Solved [154]
Theorem 5.2
$x^{2p}+by^{2}=z^{5}$ $p\geq 7$, $b$ positive odd, Partial [152]
$5\nmid h(-b)$
@@
$x^{2}+2^{\alpha}5^{\beta}11^{\gamma}y^{2p}=z^{5}$ $\alpha\geq 3$, $p>11$, Solved [152]
$\beta\geq 1$, $\gamma\geq 0$
$x^{2}+2^{\alpha}5^{\beta}19^{\gamma}y^{2p}=z^{5}$ $\alpha\geq 2$, $p\geq 11$, $p\neq 19$, Solved [152]
$\beta\geq 1$, $\gamma\geq 0$
@@
$p\equiv 47,65,113,$
139,143 or 167 (mod 168)
$ax^{2}+y^{2n}=4z^{3}$ $a\in\{11,19,43,67,163\}$, Solved [38]
$n\geq 2$
$x^{2}+ay^{2n}=4z^{3}$ $a\in\{11,19,43,67,163\}$, Solved [38]
$n\geq 2$